OpenStudy (anonymous):

How am I supposed to do this!!?? F=f o g (f times g) H(x)=x(x-2f(x))^4 g(3)=-2 g'(3)=4 f(3)=2 f'(3)=1 f'(-2)=-1 Find F'(3) and H'(3)???

OpenStudy (jamesj):

The product rule of differentiation. What is the derivative of fg, just in general terms?

OpenStudy (anonymous):

f(x) times g'(x) + g(x) times f'(x)?

OpenStudy (jamesj):

Yes, so use that for F = fg

OpenStudy (anonymous):

ok hold on lemme see if i get it

OpenStudy (anonymous):

i got the wrong answer...what am i doing wrong? i did 2(4)+-2(1)?

OpenStudy (anonymous):

well the way i read your problem, it is not f times g it looks like \[F=f\circ g\] not \[F=f\times g\]

OpenStudy (anonymous):

yea its the first one with the circle

OpenStudy (anonymous):

so you need the "chain rule" rather then the product rule

OpenStudy (jamesj):

In which case you need the chain rule

OpenStudy (anonymous):

i forget the chain rule...

OpenStudy (anonymous):

\[(f\circ g)'=f'(g)\times g'\]

OpenStudy (anonymous):

or if you prefer \[(f\circ g)'(x)=f'(g(x))\times g'(x)\] try those numbers for \[(f\circ g)'(2)=f'(g(2))\times g'(2)\]

OpenStudy (anonymous):

shouldnt i be doing it for 3not 2?

OpenStudy (anonymous):

oh sorry it is \[(f\circ g)'(3)=f'(g(3))\times g'(3)\]

OpenStudy (anonymous):

yes my mistake

OpenStudy (anonymous):

haha okay so how would i do g(3)=-2 inside of f'(3)=1? 1 times -2?

OpenStudy (jamesj):

(I should have seen that this was the composite, thanks sat73)

OpenStudy (anonymous):

how do i do f'(g(3))?

OpenStudy (jamesj):

Well, g(3) = -2, hence f'(g(3)) = f'(-2) = ...

OpenStudy (anonymous):

ohh okayy